# Brain Foundation Models: A Survey on Advancements in Neural Signal Processing and Brain Discovery

### v1

@@ II-C2 Training Strategy
Remark: Although SSL has become the dominant training paradigm for BFMs, its suitability for brain data remains uncertain. In NLP and CV, SSL benefits from structured data and rich semantic information—for instance, BERT [7] captures word co-occurrence patterns, while CLIP [24] aligns images with text. However, brain signals exhibit high noise, low reproducibility, and substantial individual variability, making traditional SSL challenging for neural signal modeling. First, neural activity patterns are unstable, varying significantly even for the same task across individuals, making it difficult to establish a consistent representation space. Second, EEG and fMRI lack hierarchical structures, meaning masked reconstruction and contrastive learning may struggle to capture generalizable features. Consequently, while SSL is highly effective in other AI fields, its practical benefits for BFMs remain uncertain and require further validation.
Furthermore, most BFMs rely heavily on fine-tuning, failing to achieve true cross-task generalization. The fundamental goal of SSL is to develop universal neural representations that can be applied across tasks without large-scale labeled data. However, nearly all existing BFMs require extensive fine-tuning for competitive performance. Notably, NeuroLM, the only BFM without fine-tuning, adopts multi-task instruction tuning to improve cross-task adaptability, yet its performance remains significantly below fine-tuned models. This suggests that while SSL offers a unified pre-training framework, current BFMs still lack zero-shot generalization, making them less versatile compared to large-scale NLP models like GPT.
Additionally, BFM training strategies are adapted mainly from NLP and CV, lacking neuroscience-specific optimizations. Techniques like masked prediction and contrastive learning fail to account for functional connectivity, hierarchical structures, and biological constraints inherent in brain activity. For example, BERT leverages word order and syntax for masked token recovery, whereas EEG signals rely on complex temporal dependencies, making naive masking potentially ineffective. Similarly, in contrastive learning, CV models assume samples of the same category share similar representations, yet in brain data, even individuals performing the same task may exhibit highly divergent neural activity patterns. As a result, BFMs remain direct adaptations of AI techniques rather than models tailored for neural signals, ultimately limiting their biological interpretability and real-world applicability.
II-D Large Scale Data and Multi-Modal Fusion
In BFMs, one of the central challenges is managing the diversity and complexity of data used in different tasks. Neural data, such as EEG, fMRI, magnetoencephalography (MEG), and functional near infrared spectroscopy (fNIRS), come in various formats, offer distinct insights into brain activity, and often vary significantly in terms of data characteristics. EEG signals capture high-frequency temporal dynamics, which is useful for the analysis of real-time cognitive processes, while fMRI offers rich spatial information about brain structure and function. These data modalities, along with other neuroimaging techniques, each represent different facets of brain activity, creating both opportunities and challenges for data integration in BFMs.
The diversity of data sources is further compounded by the varying experimental conditions under which they are collected. In tasks like brain disease diagnosis, neurorehabilitation, or cognitive state monitoring, the model may encounter data from different individuals, hardware, and environmental settings. Each of these factors introduces variability, such as individual brain anatomy differences in fMRI or inter-subject variances in EEG signal patterns. For the model to generalize effectively across these variations, it must be capable of learning from diverse and heterogeneous data in a manner that maintains task-specific relevance while minimizing biases caused by data inconsistencies.
Thus, BFMs must be robust enough to handle large-scale, multi-source datasets spanning diverse brain signals and experimental conditions, facilitating the discovery of generalizable patterns while accommodating individual differences in neural activity. This ability to process diverse datasets efficiently is essential for the model’s success in a wide range of neuroscience applications, from diagnosing brain disorders to developing personalized cognitive enhancement strategies.
Given the multimodal nature of brain data, multi-modal fusion is a fundamental component in the design and training of BFMs. Integrating multiple data types, such as EEG and fMRI, allows the model to leverage the complementary strengths of each modality. However, the fusion process introduces significant challenges, particularly due to the differences in temporal and spatial resolution, as well as variations in data formats and dimensionalities across modalities.
In multimodal fusion, the goal is to integrate data from various modalities so the model can leverage the unique information offered by each source. For example, EEG provides high temporal resolution, capturing brain activity fluctuations over time, while fMRI offers spatially detailed maps of brain activity. To effectively utilize both, BFMs must adopt fusion techniques that can synchronize and combine these diverse signals into a cohesive representation that is informative across multiple tasks.
@@ III-A Brain Decoding
Brain decoding refers to the use of BFMs to interpret neural signals and decode mental states, intentions, or cognitive processes from brain activity. This process is critical for various applications, from BCIs to brain disease diagnostics. BFMs are particularly valuable in this domain due to their ability to generalize across diverse tasks, modalities, and scenarios, allowing for more robust and adaptable models for interpreting neural signals. Their flexibility makes them essential for advancing brain research and improving clinical and real-world applications. The typical BFMs for brain decoding are summarized in Tab. II.
@@
Model Pretraining Fine-tuning Cross Scenarios Cross Tasks Multi Modalities
BrainLM [19] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
LaBraM [25] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
NeuroLM [20] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Brant [14] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
BrainSegFounder [26] $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
@@
BRAINBERT [29] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
@@
Brain-JEPA [13] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
FoME [30] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
@@
Neuro-GPT [34] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$
@@
CEReBrO [37] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
CBraMod [16] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
EEGPT-Yue [38] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
TGBD [23] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
@@
EEGPT [17] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
Brant-X [22] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
FM-APP [39] $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$
@@ III-A1 Pretraining Only
The Pretraining Only approach trains BFMs on large, diverse datasets of neural signals without specific task adaptation. The model learns universal representations that capture core patterns of brain activity shared across various contexts, scenarios, and experimental conditions. This makes it highly applicable in cross-scenario tasks, where neural data may vary across a wide range of conditions, such as:
  • Epilepsy Detection Across Multiple Scenarios: In epilepsy diagnosis, BFMs pretrained on EEG data from diverse seizure scenarios, such as different seizure types, environmental conditions, and patient populations, can generalize across different seizure manifestations. By leveraging these generalized features, the pretrained model can effectively recognize seizures in new datasets that may differ in subjects, recording settings, or seizure types. This capability is crucial for developing real-time seizure detection systems for clinical use.
  • Wide Range Speech-BCI Data Decoding: For speech-evoked BCI systems, BFMs pretrained on diverse speech-related brain data captured from different subjects, speech tasks, and sensory modalities can decode speech intentions more effectively across multiple scenarios. Whether the data comes from speech imagery, silent speech, or actual speech, the pretrained model can handle this variability and decode speech-related mental states. This application is critical for improving the accessibility of communication devices for individuals with speech impairments.
@@ III-A2 Pretraining + Fine-tuning
The Pretraining + Fine-tuning approach combines the power of pretrained models with task-specific adaptation. After pretraining on large-scale datasets, BFMs are fine-tuned for specific tasks, enhancing their performance in specialized applications. This approach is particularly effective in cross-task scenarios, where the same model needs to be adapted for multiple different types of BCI tasks. This flexibility is advantageous for applications that require the model to work across a range of BCI paradigms, such as:
  • Cross-Task BCI Applications: A model pretrained on diverse neural data can be fine-tuned for various BCI tasks. For example, a single model can be adapted for motor imagery, cognitive workload assessment, and emotion recognition. This enables a unified framework capable of handling multiple BCI applications without task-specific model design, making the model more versatile and reducing development time and costs.
  • Generalist Disease Diagnosis: Pretraining on general neural signal data, followed by fine-tuning for specific diseases (e.g., PTSD, anxiety, and neuroticism), enables BFMs to identify biomarkers for disease diagnosis and progression. This approach can be particularly useful for developing diagnostic tools that can work across a wide range of patients, reducing the need for collecting large amounts of task-specific training data. By leveraging the generalized knowledge from pretraining, the fine-tuned model can provide accurate and reliable predictions for various diseases based on neural data.
@@ III-B1 Pretraining + Interpretability
In the Pretraining + Interpretability approach, BFMs are pretrained on large-scale neural signal datasets, such as EEG or fMRI, to learn general brain activity patterns. These pretrained models are then used to construct a digital brain by simulating biological brain processes. Through the use of interpretability techniques, such as perturbation analysis or attention mechanisms, researchers can analyze the digital brain and gain insights into how the biological brain operates under various conditions. This approach is particularly valuable in exploring the complex dynamics of the human brain and uncovering mechanisms that are difficult to study directly. Key applications include:
  • Digital Brain Construction: Using BFMs, researchers can build a digital brain [40], a computational model that mimics the biological brain’s functioning. This model can simulate neural processes across various brain regions and networks. By analyzing it, researchers can explore how neural circuits operate, how different regions communicate with each other, and how the brain processes information during cognitive tasks. These simulations enable the discovery of new neural mechanisms that may not be observable through direct measurements alone.
  • Exploring Brain Function and Disease Mechanisms: The digital brain provides a platform for studying brain functions in both healthy and diseased states. For instance, perturbation analysis can simulate disruptions in brain activity, such as those caused by neurodegenerative diseases like Alzheimer’s or Parkinson’s. By analyzing how these disruptions affect neural networks in the digital brain, researchers can gain deeper insights into disease progression and identify potential biomarkers or therapeutic targets for early intervention.
  • Understanding Cognitive Processes: Through the use of digital brain models, BFMs can be used to simulate complex cognitive processes, such as memory, decision-making, and learning. Researchers can explore how different brain regions and networks interact to perform these tasks, revealing the underlying neural mechanisms of cognition. This approach could also help identify the brain regions involved in specific cognitive disorders, leading to better diagnostic tools and more targeted treatments.
  • Uncovering Unknown Mechanisms in Biological Brain: By studying the digital brain, researchers can uncover new insights into brain function that are difficult to observe directly in biological systems. For example, interpretability methods can be applied to identify which brain regions or networks play critical roles in various cognitive functions or emotional responses. These discoveries could reveal unknown neural mechanisms that contribute to behavior and cognition, providing new avenues for research into mental health, learning processes, and brain plasticity.
@@ IV Future Directions and Challenges
IV-A Data for BFMs
One of the most critical aspects of BFMs is data availability, diversity, and quality. Future research should focus on incorporating a wide range of brain data types, including fMRI for capturing dynamic brain activity, EEG and magnetoencephalography (MEG) for high-temporal-resolution neural signals, diffusion tensor imaging (DTI) for mapping white matter connectivity, and structural MRI for anatomical insights. Additionally, integrating genetic information, clinical reports, and behavioral assessments can enhance the model’s ability to understand neurological disorders and cognitive functions holistically. To overcome the challenge of limited data, generative models (e.g., GANs and diffusion models) can be used to synthesize realistic brain data, filling gaps in underrepresented populations or rare disease cases. Furthermore, privacy concerns hinder large-scale data sharing, making federated learning a promising approach to enable decentralized training across institutions while maintaining patient confidentiality.
However, several challenges remain: (1) Data scarcity and inconsistency due to variations in imaging protocols, scanner settings, and subject demographics, leading to high variability across datasets and limiting model generalization. (2) The lack of standardized preprocessing pipelines, making it difficult to harmonize data from different sources and reducing the reproducibility of BFM training. (3) Ethical and legal constraints, such as GDPR, HIPAA, and institutional policies, which restrict access to brain data and hinder large-scale collaborative learning.
IV-B Architectural Innovations and Backbone Selection
The choice of backbone architecture is crucial for capturing spatial-temporal patterns in brain data, as different brain modalities present unique challenges in representation learning. CNNs are effective for processing spatial patterns in structural MRI and fMRI but struggle with long-range dependencies. Graph Neural Networks (GNNs) leverage brain connectivity structures, making them well-suited for modeling functional and structural networks. LSTMs and Transformers excel in handling temporal dependencies in EEG and MEG data, where sequential information is key, but their applicability to volumetric brain data remains underexplored. Future research should focus on developing hybrid architectures that combine the strengths of these models while incorporating biological priors, brain connectivity, and functional networks.
However, several challenges remain: (1) No single architecture is universally optimized for all brain modalities, as EEG and fMRI have vastly different data properties, requiring specialized processing techniques; (2) The computational complexity of Transformers and large-scale models makes training and deployment resource-intensive, necessitating advancements in model compression and pruning to improve scalability; and (3) Incorporating neuroscientific domain knowledge into deep learning architectures remains difficult, as traditional AI models are often purely data-driven and lack explicit integration of neurobiological constraints, such as hierarchical brain organization or causal relationships in neural activity.
IV-C Training Strategies and Efficiency
Training BFMs efficiently requires novel learning paradigms that can leverage the large amount of unlabelled brain data while ensuring biological consistency. SSL can help extract meaningful representations without relying on expensive expert annotations, while contrastive learning can enhance feature discrimination by learning similarities and differences between brain states. Meta-learning and transfer learning may also improve generalization by enabling BFMs to adapt quickly to new datasets or tasks with minimal retraining. A critical direction for future research is developing multimodal training strategies that integrate diverse data types, such as EEG (high temporal resolution), fMRI (spatial brain activity), MEG, MRI, and genetic data, into a unified framework. Additionally, multimodal fusion requires handling differences in spatial and temporal resolution, as well as aligning heterogeneous data distributions. To make BFMs scalable and accessible, reducing training costs is crucial. Techniques such as knowledge distillation, model quantization, and efficient hardware utilization (e.g., using TPUs or distributed computing) can help optimize resource use.
However, several challenges persist: (1) Limited availability of high-quality labeled data, training BFMs highly relies on SSL, which needs meticulous designs; (2) Complexity in fusing multimodal data, as EEG and MRI provide fundamentally different types of information; and (3) High computational costs, requiring efficient training frameworks that balance accuracy, scalability, and accessibility for broader research adoption.
IV-D Usage of The Pretrained BFMs
BFMs have the potential to transform neuroscience and medicine by enabling early disease detection, cognitive assessment, and BCIs. Future research should focus on improving fine-tuning techniques to efficiently adapt pretrained models for specific clinical and cognitive tasks, minimizing the need for large labelled datasets. Additionally, integrating LLMs with BFMs could create interactive neuro-AI systems capable of interpreting complex brain data, assisting researchers and clinicians in diagnostics, therapy planning, and personalized medicine.
However, several challenges must be addressed: (1) Domain adaptation and generalization, as BFMs must perform reliably across different populations, imaging modalities, and tasks; (2) Inference speed limitations, especially for real-time BCIs that require ultra-low latency for neurofeedback applications; and (3) Trustworthiness and explainability, as medical adoption depends on transparent, interpretable AI decisions that align with established neuroscience principles.
IV-E Interpretability and Explainability
Interpreting BFMs is crucial for clinical trust, neuroscientific insight, and ethical deployment in real-world applications. Future research should focus on explainable AI (XAI) techniques to make model decisions more transparent and interpretable. Methods such as attention maps, saliency techniques, and concept-based interpretability can help identify which brain regions contribute to specific predictions. Additionally, neuro-symbolic AI, which combines deep learning with neuroscientific rules, may improve interpretability by aligning model outputs with known brain functions. Integrating multimodal large language models (MLLMs) with BFMs presents another promising direction. MLLMs can generate textual explanations for model predictions, helping neuroscientists and clinicians understand brain patterns more intuitively. Visualization tools that link brain activity patterns with natural language descriptions may enhance human-AI collaboration in research and clinical settings.
However, several challenges remain: (1) The black-box nature of deep learning models makes it difficult to explain how BFMs derive their predictions; (2) The lack of standardized interpretability metrics for brain AI models hinders reliable assessment across studies; and (3) Potential for misleading explanations, where AI-generated insights may not align with established neuroscience principles, necessitating careful validation.
IV-F Ethical, Regulatory, and Standardization Considerations
As BFMs move toward real-world deployment, ensuring ethical AI practices, fairness, and regulatory compliance is critical for their acceptance in neuroscience and healthcare. Future research should focus on developing standardized benchmarks and evaluation frameworks to ensure transparency, reproducibility, and robustness across different datasets, populations, and clinical applications. Establishing clear validation protocols will be essential for gaining trust among medical professionals and regulatory bodies.
However, several challenges remain: (1) Regulatory hurdles, such as compliance with FDA, GDPR, and HIPAA, which impose strict data privacy and safety requirements; (2) Potential biases in brain models, as training data may underrepresent certain demographics, leading to disparities in clinical decision-making; and (3) The lack of standardized evaluation frameworks, making it difficult to assess model performance across institutions and delaying real-world adoption.

### v2

@@ II-C2 Training Strategy
Remark: Although brain-heterogeneous modality contrastive learning currently offers limited improvements for direct brain data analysis, its long-term value cannot be overlooked. As model capabilities expand, the next phase of BFMs will place greater emphasis on flexible input-output modes to accommodate diverse brain data formats and task requirements. This strategy enables BFM to gradually align with current mainstream multimodal large language models, thereby acquiring the ability to handle complex semantic alignment and cross-modal understanding, providing the necessary foundational support for more advanced BCI tasks.
Fine-tuning & Task-Specific Adaptation: Once the BFM has been pre-trained using SSL and masking techniques, the model proceeds to the fine-tuning phase, where it is adapted for specific tasks using labeled data. The fine-tuning process involves supervised learning, where the model is trained to predict specific outcomes (e.g., brain disease diagnosis or cognitive state classification) based on the representations learned during pre-training.
Fine-tuning typically uses cross-entropy loss for classification tasks or mean squared error for regression tasks. Let $D=\left\{\left(x_{i},y_{i}\right)\right\}$ represent the labeled dataset, where $x_{i}$ is the input brain signal, and $y_{i}$ is the corresponding label. The fine-tuning objective is to minimize the task-specific loss:
$\mathcal{L}_{\text{fine-tune }}=\sum_{i}\mathcal{L}_{\text{task }}\left(f\left(x_{i}\right),y_{i}\right)$ (6)
For classification tasks, this might involve predicting discrete classes of brain activity or disease, while for regression tasks, the output might represent continuous values, such as cognitive scores or clinical measures.
Remark: Current exploration of fine-tuning in BFMs remains insufficient. Compared to text and images, brain data exhibits characteristics such as significant individual variability and unstable cross-task distribution, making BFMs more reliant on fine-tuning to achieve effective transfer. However, strategies such as unsupervised domain adaptation, test-time training, and efficient parameter fine-tuning methods (e.g., Adapter, LoRA) have not yet been widely applied in BFMs. By drawing on these mechanisms that have been validated as effective in NLP and CV, it is likely that the adaptability and practicality of BFMs in real-world BCI scenarios can be significantly enhanced.
Summary of Training Strategy: The suitability of SSL for brain data is uncertain due to high noise, variability, unlike NLP and CV data. These challenges hinder learning generalizable features with standard SSL, leading most current BFMs to rely heavily on extensive fine-tuning. This reliance curtails true cross-task generalization and zero-shot capabilities, even for models like NeuroLM. Furthermore, many BFMs training strategies are merely adaptations from NLP and CV, lacking neuroscience-specific optimizations. Consequently, masked prediction or contrastive learning often fails to incorporate vital biological constraints, functional connectivity, or complex brain dynamics. This positions current BFMs more as direct AI adaptations than neuro-tailored models, limiting their biological interpretability and real-world applicability.
@@ III-A Brain Decoding
Brain decoding refers to using BFMs to interpret neural signals and decode mental states, intentions, or cognitive processes from brain activity. This process is critical for various applications, from BCIs to brain disease diagnostics. BFMs are particularly valuable in this domain due to their ability to generalize across diverse tasks, modalities, and scenarios, allowing for more robust and adaptable models for interpreting neural signals. Their flexibility makes them essential for advancing brain research and improving clinical and real-world applications. The typical BFMs for brain decoding are summarized in Tab. II and Table III, which provide a quantitative performance comparison of representative models of fMRI and EEG data on public datasets, respectively.
@@
Model Pretraining Fine-tuning Cross Scenarios Cross Tasks Multi Modalities Size
BrainLM [21] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 650M
LaBraM [26] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 369M
NeuroLM [19] $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 1.7B
Brant [15] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 505M
BrainSegFounder [27] $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ 69M
@@
BRAINBERT [30] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 43M
@@
Brain-JEPA [14] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ 307M
FoME [31] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 745M
@@
Neuro-GPT [35] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ 80M
@@
CEReBrO [38] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ 85M
CBraMod [17] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 4M
EEGPT-Yue [39] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ 1.1B
TGBD [25] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$
@@
EEGPT [18] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ 10M
Brant-X [24] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ 1B
FM-APP [40] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
@@ III-A1 Pretraining Only
The Pretraining Only approach trains BFMs on large, diverse datasets of neural signals without specific task adaptation. The model learns universal representations that capture core patterns of brain activity shared across various contexts, scenarios, and experimental conditions. This makes it highly applicable in cross-scenario tasks, where neural data may vary across a wide range of conditions, such as: 1) Epilepsy Detection Across Multiple Scenarios: In epilepsy diagnosis, BFMs pretrained on EEG data from diverse seizure scenarios, such as different seizure types, environmental conditions, and patient populations, can generalize across different seizure manifestations. By leveraging these generalized features, the pretrained model can effectively recognize seizures in new datasets that may differ in subjects, recording settings, or seizure types. This capability is crucial for developing real-time seizure detection systems for clinical use. 2) Wide Range Speech-BCI Data Decoding: For speech-evoked BCI systems, BFMs pretrained on diverse speech-related brain data captured from different subjects, speech tasks, and sensory modalities can decode speech intentions more effectively across multiple scenarios. Whether the data comes from speech imagery, silent speech, or actual speech, the pretrained model can handle this variability and decode speech-related mental states. This application is critical for improving the accessibility of communication devices for individuals with speech impairments.
@@ III-A2 Pretraining + Fine-tuning
The Pretraining + Fine-tuning approach combines the power of pretrained models with task-specific adaptation. After pretraining on large-scale datasets, BFMs are fine-tuned for specific tasks, enhancing their performance in specialized applications. This approach is particularly effective in cross-task scenarios, where the same model needs to be adapted for multiple different types of BCI tasks. This flexibility is advantageous for applications that require the model to work across a range of BCI paradigms, such as: 1) Cross-Task BCI Applications: A model pretrained on diverse neural data can be fine-tuned for various BCI tasks. For example, a single model can be adapted for motor imagery, cognitive workload assessment, and emotion recognition. This enables a unified framework capable of handling multiple BCI applications without task-specific model design, making the model more versatile and reducing development time and costs. 2) Generalist Disease Diagnosis: Pretraining on general neural signal data, followed by fine-tuning for specific diseases (e.g., PTSD, anxiety, and neuroticism), enables BFMs to identify biomarkers for disease diagnosis and progression. This approach can be particularly useful for developing diagnostic tools that can work across a wide range of patients, reducing the need for collecting large amounts of task-specific training data. By leveraging the generalized knowledge from pretraining, the fine-tuned model can provide accurate and reliable predictions for various diseases based on neural data.
@@
TABLE III: Comparison of Typical fMRI and EEG based BFM models. MSE (Mean Squared Error), $\rho$ (Pearson Correlation Coefficient), ACC (Accuracy), F1 (F1 Score), AUROC (Area Under the Receiver Operating Characteristic Curve).
Dataset Task Metrics BrainLM[21] Brain-JEPA[14] NeuroLM [19] LaBraM[16]
fMRI-based Models
HCP-Aging Age MSE $0.331\pm 0.018$ $0.298\pm 0.017$
Age $\rho$ $0.832\pm 0.028$ $0.844\pm 0.030$
Sex ACC $74.39\pm 1.55$ $81.52\pm 1.03$
Sex F1 $77.51\pm 1.13$ $84.26\pm 0.82$
Neuroticism MSE $0.942\pm 0.082$ $0.897\pm 0.055$
Neuroticism $\rho$ $0.231\pm 0.012$ $0.307\pm 0.006$
Flanker MSE $0.971\pm 0.054$ $0.972\pm 0.038$
Flanker $\rho$ $0.318\pm 0.048$ $0.406\pm 0.027$
ADNI NC/MCI ACC $75.79\pm 1.05$ $76.84\pm 1.05$
NC/MCI F1 $85.66\pm 1.27$ $86.32\pm 0.54$
[…]
EEG-based Models
TUAB Abnormal Detection ACC $0.7969\pm 0.0091$ $0.8140\pm 0.0019$
Abnormal Detection AUROC $0.7884\pm 0.0194$ $0.9022\pm 0.0009$
TUEV Event Type Classification ACC $0.4679\pm 0.0356$ $0.6409\pm 0.0065$
Event Type Classification F1 Score $0.7359\pm 0.0219$ $0.8312\pm 0.0052$
SEED Emotion Recognition ACC $0.6034\pm 0.0010$ $0.7318\pm 0.0019$
Emotion Recognition F1 Score $0.6063\pm 0.0030$ $0.7354\pm 0.0021$
@@ III-B1 Pretraining + Interpretability
In the Pretraining + Interpretability approach, BFMs are pretrained on large-scale neural signal datasets (e.g., EEG, fMRI) to learn general patterns of brain activity. These models are then fine-tuned and structurally constrained using multimodal brain atlases (e.g., Brainnetome) to construct a Digital Twin Brain (DTB) [41]. The modeling pipeline integrates a multiscale architecture: the microscopic level models individual neurons with detailed electrical dynamics (e.g., Hodgkin–Huxley, integrate-and-fire); the mesoscopic level captures interactions within and between excitatory and inhibitory populations (e.g., Wilson-Cowan, dynamic mean-field); and the macroscopic level simulates large-scale activity and inter-regional functional connectivity using whole-brain network models. Training aims to reproduce empirical functional connectivity and signal variability via model inversion or parameter exploration. Validation is performed by comparing simulated outputs with neural data and assessing responses to perturbation-based simulations.
Through interpretability techniques such as perturbation analysis or attention mechanisms, researchers can analyze the digital brain and gain insights into how the biological brain operates under various conditions. This approach is particularly valuable for exploring complex neural dynamics and uncovering mechanisms that are difficult to observe directly. Key applications include: 1) Exploring Brain Function and Disease Mechanisms: The digital brain serves as a platform for studying brain function in both healthy and diseased states. Perturbation analysis can simulate disruptions such as those from neurodegenerative diseases (e.g., Alzheimer’s, Parkinson’s), allowing researchers to examine network-level effects, understand disease progression, and identify potential biomarkers or therapeutic targets. 2) Understanding Cognitive Processes: Digital brain models enable simulation of complex cognitive functions like memory, decision-making, and learning. By examining interactions among brain regions, researchers can uncover neural mechanisms underlying cognition and identify regions linked to cognitive disorders, aiding diagnosis and targeted treatment. 3) Uncovering Unknown Mechanisms in the Biological Brain: DTB models offer access to neural processes difficult to observe in vivo. Interpretability methods help reveal which regions or networks are critical for specific cognitive or emotional functions, offering new insights into behavior, mental health, and brain plasticity.
@@ IV Future Directions and Challenges
IV-A Data for Brain Foundation Models
Data availability, diversity, and quality are foundational to the success of BFMs. Future research should prioritize the integration of diverse brain imaging modalities such as fMRI for dynamic activity patterns, EEG and magnetoencephalography (MEG) for high temporal resolution, diffusion tensor imaging (DTI) for structural connectivity, and structural MRI for anatomical mapping. Incorporating behavioral assessments, clinical labels, and genetic information will further enable more holistic modeling of brain function and disorders. To address limited data coverage across populations and conditions, generative approaches such as GANs or diffusion models can be explored to synthesize realistic brain signals, particularly for rare diseases and underrepresented cohorts.
However, key challenges remain. The variability in imaging protocols, scanner configurations, and population demographics leads to inconsistent datasets that reduce model generalizability. The absence of standardized preprocessing and annotation pipelines further complicates data harmonization and reproducibility. Addressing these issues requires the development of community-adopted protocols for cross-site harmonization, along with publicly available benchmarks for evaluating BFM performance. Moreover, technical strategies to meet data protection regulations such as GDPR and HIPAA must be tightly integrated into model design and deployment.
IV-B Architectural Innovations and Neuroscience Integration
BFMs face unique architectural challenges due to the brain’s highly structured, multimodal, and dynamic nature. While models such as CNNs, Transformers, and RNNs offer a foundation, they must be adapted to reflect domain-specific constraints. A key research direction involves integrating neuroscientific priors directly into model design. For instance, brain connectivity matrices derived from DTI or functional connectivity analyses can be used to define the topologies of Graph Neural Networks (GNNs), encoding structural or functional constraints as part of the model architecture. This approach enables spatial relationships between brain regions to be explicitly modeled, rather than inferred indirectly.
Beyond structural priors, models should also be encouraged to reflect known functional hierarchies and spectral dynamics. One promising direction is the design of loss functions that incorporate neuroscientific principles, such as promoting representational consistency with known brain oscillations (e.g., alpha, beta, and gamma bands) or penalizing divergence from established cortical gradients. Incorporating these constraints can guide model learning toward biologically plausible representations and may improve transferability across tasks and populations.
IV-C Training Strategies and Multimodal Fusion
Training BFMs effectively requires learning paradigms that can extract meaningful patterns from vast quantities of unlabeled data while respecting the constraints of biological realism. SSL and contrastive learning methods are particularly suitable, as they reduce dependence on manual annotations while capturing rich neural dynamics. These methods must be adapted to account for domain-specific objectives, such as predicting inter-regional correlations or simulating signal transitions across cognitive states. Transfer learning and meta-learning strategies can further help BFMs generalize across different populations and tasks with limited fine-tuning.
A major research direction involves the development of unified training frameworks that integrate data from multiple modalities, such as EEG, fNIRS, fMRI, MEG, and even genetics, into a coherent representational space. Aligning these modalities poses technical challenges due to their differences in spatial and temporal resolution, data scale, and noise characteristics. Advances in multimodal fusion techniques, such as cross-modal encoders or shared latent spaces, are needed to enable effective integration. Furthermore, training costs must be addressed through hardware-efficient strategies, such as distributed computing and lightweight architectures, to ensure the scalability and accessibility of BFMs for both academic and clinical settings.
IV-D Adaptation and Use of Pretrained BFMs
Once pretrained, BFMs offer significant potential for downstream applications, including early disease diagnosis, cognitive modeling, and BCI development. Fine-tuning these models for specific clinical or research tasks requires protocols that are both data-efficient and modality-aware. Emerging approaches such as prompt tuning or adapter modules could provide flexible, parameter-efficient strategies for adaptation without extensive retraining.
In parallel, integrating BFMs with LLMs opens new opportunities for developing interpretable, interactive neuro-AI systems. These systems could interpret complex brain signals and provide natural language explanations for predictions, thereby supporting clinical decision-making or cognitive assessments. Nonetheless, several challenges need to be resolved, including ensuring robust domain adaptation across datasets and populations, maintaining low-latency inference for real-time applications like BCIs, and ensuring the trustworthiness of model outputs through validation mechanisms grounded in neuroscientific principles.
IV-E Interpretability and Evaluation Standards
Interpretability is a critical requirement for the acceptance and safe deployment of BFMs in neuroscience and clinical contexts. Traditional deep learning models often act as black boxes, making it difficult to understand or validate their decisions. Future work must focus on designing interpretability techniques that are both technically rigorous and neuroscientifically meaningful. This includes developing attention mechanisms and saliency-based visualizations that highlight relevant brain regions, as well as concept-based approaches that map predictions to known cognitive or clinical constructs.
To move beyond qualitative insights, the field needs standardized interpretability metrics specifically tailored to brain modeling. These could include overlap with task-evoked activation maps, correspondence with known anatomical regions, or alignment with expert annotations. Neuro-symbolic models that combine statistical learning with interpretable rules derived from brain science offer another promising direction. Moreover, combining BFMs with multimodal language models could help translate activation patterns into human-readable explanations, fostering stronger collaboration between AI systems and clinicians or neuroscientists.